The light shining on a diffraction grating has a wavelength of 481 nm (in vacuum). The grating produces a second-order bright fringe whose position is defined by an angle of 8.59°. How many lines per centimeter does the grating have?
Answer #1

Solution)

We know,

m λ = d sinθ   

So,

d = m λ / sinθ = 2 * 481 nm /sin 8.59 = 6440.69nm = 0.0006440 cm perline

Now,

The number of lines per cm is the inverse -

Hence,

1 / 0.0006440 cm per line =1552.79 lines per cm

==============